OpenStudy (elenathehomeschooler):

1. For each of the following functions, (a) describe the domain and range, and (b) sketch the graph. a. f(x) ^5sqr(x-4)+2

OpenStudy (elenathehomeschooler):

@jim_thompson5910 can you help me ring the range and domain?

OpenStudy (elenathehomeschooler):

*find

OpenStudy (anonymous):

is it \[f(x)=\sqrt[5]{x-4}+2\]?

OpenStudy (elenathehomeschooler):

yes

OpenStudy (anonymous):

you can take an odd root (3,5,7,...) of ANY number

OpenStudy (anonymous):

i.e. the there are no restrictions on the domain

OpenStudy (elenathehomeschooler):

so the domain would be all real numbers?

OpenStudy (anonymous):

yes

OpenStudy (elenathehomeschooler):

how do i find the range?

OpenStudy (anonymous):

as x gets larger and larger, so does \(\sqrt[5]{x}\) i.e as \(x\to \infty\) your function goes to infinity as well

OpenStudy (anonymous):

and as \(x\to -\infty\) so does your function

OpenStudy (anonymous):

that is not really an explanation, though, that is just a fact not sure how to explain it in words

OpenStudy (anonymous):

and odd root will have this same property

OpenStudy (elenathehomeschooler):

so for range it would be \[\infty,-\]

OpenStudy (anonymous):

all real numbers just like the domain

OpenStudy (elenathehomeschooler):

\[(\infty,-\infty)\]

OpenStudy (anonymous):

oh no you don't put the \(-\infty\) on the right! it goes on the left

OpenStudy (elenathehomeschooler):

oh okay